# Learning Ising Models under Hard Constraints using One Sample

### v1

@@ 1 Introduction
Markov random fields (MRFs) are a common framework for analyzing high-dimensional distributions with complex conditional independence structures. A well-studied example of an MRF and the primary topic of inquiry in this paper is the Ising model Isi (25), a probability measure $\mu_{G,\beta}$, over all assignments $\bm{\sigma}$ in the binary hypercube $\{-1,1\}^{n}$. The model is parameterized by a graph $G=(V,E)$ and inverse temperature $\beta$, and taking the form $\mu_{G,\beta}\propto\exp\left(\beta\bm{\sigma}^{\top}A\bm{\sigma}\right)$, with $A$ being the adjacency matrix of $G$. The simplicity of the Ising model has led to widespread adoption in fields as disparate as statistical physics, finance, the social sciences, and computer vision, among others (see Cha (07); QZ (11); Har (13); WCMN (15) and the references therein for a brief collection of examples). These applications have, in turn, motivated a substantial body of research on efficient sampling Bre (15); LS (13); SS (12), rigorous testing DDK (19), and principled inference of the inverse temperature parameter and interaction matrix DDDK (20); DDD+ (21) under the framework of the Ising model.
Beginning from the work of Cha (07), there has been substantial interest in the task of estimating inverse temperature parameter $\beta$, given only one sample $\bm{\sigma}\sim\mu_{G,\beta}$ and the graph $G$, using the maximum pseudolikelihood estimator of Bes (75). This setup subsumes the setting of multi-sample estimation as $\ell$ samples of an Ising model over $n$ nodes is equivalently a single sample of an $n\ell$ Ising model over $\ell$ disconnected components of a graph. This line of work is driven by the inherent technical constraints of network data, wherein it is often impractical to obtain independent observations of the same network responses DDP (19); DDDK (20). Interestingly, despite both static and dynamic phase transitions in model behavior as $\beta$ varies which can render sampling computationally intractable (NP-hard) GSV (16), it remains possible to construct consistent estimators of $\beta$ and the interaction matrix $A$, provided that $\beta=\mathcal{O}(1)$ DDDK (20); DDD+ (21); DDP (19); MNH+ (22).
@@
In many real-world applications, however, we face not only soft constraints, which influence the model behavior by introducing correlations or dependencies, while the full support of the measure remains intact, but also hard constraints: Certain configurations are outright forbidden, and entire regions of the configuration space are excluded from the support of the distribution. Such hard-constrained models (also called truncated) arise naturally in applications involving high-dimensional, interconnected systems with strict feasibility requirements, one particularly notable example of which arises in the context of spatial transcriptomics—a biological framework for characterizing gene and protein expression in cells within organic tissue, relative to their spatial organization. The local relationships between cells are often represented as nodes in a graph with edges linking cells together that are close in physical space (ELZ+, 19). Associated with each node is a set of measurements of expression to capture how the relationships between cells impact the phenotype of a given node. The complex relationship between genes often forces certain configurations of expressions to be infeasible. In fact, the phenomenon of lateral inhibition can cause a cell expressing gene one to prevent its neighbors from expressing it as well, instead causing them to express gene two, as seen in the Notch-Delta pathway GT (01). The hard constraints discussed in the above setting are not unique to spatial transcriptomics and are also commonly found in the context of channel assignments in communication networks (ZM, 06), carrier-sense multiple access networks (DT, 06; DDT, 09), and multicasting networks (KAV, 02; LRZ, 06) among others.
@@
In this work, we study the problem of parameter estimation in $n$-dimensional Ising models that are hard-constrained to the satisfying assignments of a bounded-degree $k$-SAT formula $\Phi$ expressed in CNF (Conjunctive Normal Form), using one sample. This means that we have access to a sample from an Ising model, conditioned that it only takes values in a subset $S\subset\{\pm 1\}^{n}$ that is represented through the satisfying assignments of a $k$-SAT formula, adopting the framework from GKK (24). Learning in truncated MRFs using one or multiple samples has been studied in the context of discrete product distributions truncated by the set of satisfying assignments of $k$-SAT formulas GKK (24); GGZ (25), more generally by truncated sets with combinatorial structure FKT (22), the hard-core model, and integer valued spins constrained over proper $H$-colorings BCŠV (18); BR (21). Our key deviation from the aforementioned works stems from the fact that the Ising model is not a product distribution, and common tools used to control the concentration of measure on the hypercube do not apply. This, moreover, induces two sources of interdependence, namely from the model itself and from the structure of the truncation set. With this background in mind, we seek to address the following challenge.
@@ 1.1 Our Results
Our main contribution is an affirmative answer to the previous challenge, by providing a sufficient condition on the $k$-SAT formula that induces the truncation set, in terms of the maximum degree $\Delta$ of $G$. We begin by formally defining the class of truncated Ising measures that is the primary inquiry of this work. Given a maximum degree $\Delta$ graph with associated adjacency matrix $A$, and inverse temperature $\beta$ we define the pmf of a truncated Ising model for any $\bm{\sigma}\in\{\pm 1\}^{n}$ to be
@@
where $\mathbf{1}$ captures the indicator function and $Z_{\beta,S}$ is a renormalization term called the partition function. We focus on the case where $S$ is expressed as the set of satisfying assignments of a bounded degree $k$-SAT formula $\Phi_{n,k,d}=\Phi$ in CNF. Formally, $\Phi$ is a collection of literals and clauses $(\mathcal{V},\mathcal{C})$, in which each clause $C\in\mathcal{C}$ contains $k$ literals, and each literal $v\in\mathcal{V}$ appears in at most $d$ distinct clauses; $d$ is called the degree of the formula. Each element $\bm{\sigma}\in\{-1,1\}^{n}$ corresponds to a truth assignment for $\Phi$, where the variable $v_{i}$ is assigned to true if $\sigma_{i}=1$ and false if $\sigma_{i}=-1$. Any subset of the hypercube can be represented as the set of satisfying assignments of a $d$ degree $k$-SAT formula, provided that $d$ is sufficiently large.
@@
Notice when $\Delta$ is $\mathcal{O}(1)$, our estimate achieves $\mathcal{O}(1/\sqrt{n})$-consistency, matching the minimax rate for parameter estimation. Likewise, the restriction of $\Delta$ to be on the order of $o(n^{1/6})$ ensures $\hat{\beta}$ is a consistent estimator.
@@ 1.2 Technical Overview
Given a single-parameter exponential family like the one we focus on, a natural approach to estimating the parameter is to find the maximum likelihood estimate. However, the computational intractability of the partition function $Z_{\beta,S}$ for Ising models (see GSV (16) and the references therein) renders this approach infeasible. In light of these challenges, we utilize the maximum pseudolikelihood estimator introduced by Bes (75) and provided below.
@@
We note that the second equality holds because $\log(\cdot)$ is a monotone function. Towards demonstrating the consistency of the maximum (log)-pseudolikelihood estimate $\hat{\beta}$, we follow the first and second derivative paradigm outlined by Chatterjee Cha (07); DDP (19); GKK (24); GGZ (25), which involves showing,
@@
The first condition ensures the derivative of the log-pseudolikelihood objective with respect to the true model parameters $\beta$ divided by $n$ is close to 0, which is the value of the gradient of $\phi$ computed at the estimator, which in turn implies $\beta$ is an approximate stationary point of the objective. Moreover, by demonstrating that the second derivative of the objective $\nabla^{2}_{\beta}\phi(\beta;A,\bm{\sigma})$ is $\Omega(n/\Delta^{3})$-strongly convex with probability 1o(1)1-o(1) over a draw of the truncated Ising model, it implies that approximate stationary points of the objective are close in Euclidean distance to the optimum. Effectively, we combine these two facts to show the proximity of the optimum of the log-pseudo-likelihood objective to $\beta$.
@@
Showing both of these conditions hold simultaneously is made complex due to the highly non-uniform measure induced by conditional dependencies of both the interaction matrix $A$ and the truncation set $S$. To demonstrate the first condition, we craft upper bounds on the variance of the first derivative of $\phi$, using the technique of exchangeable pairs pioneered by Cha (07), which, when combined with Chebyshev’s inequality, implies an upper bound in probability. The primary challenge of this work lies in establishing the second condition; unlike previous works which used the deterministic structure of the interaction matrix to guarantee the concavity of the objective, in our setting the second derivative is instead determined by the number of elements Hamming distance  [^1] The Hamming distance is a metric which measures the number of indices in which two vectors differ. If a neighbor $\bm{\sigma}^{\prime}$ is Hamming distance one away from $\bm{\sigma}$, it implies we can flip one index of $\bm{\sigma}$ to yield $\bm{\sigma}^{\prime}$ one away from the sample $\bm{\sigma}$ in the truncation set, and the magnitude of the magnetizations $m_{i}(\bm{\sigma})$ at these neighbors. We begin by showing a lower bound on $m_{i}(\bm{\sigma})$ in probability via a coupling argument which exploits the underlying edge structure of the connectivity graph $G$.
@@
To show that, with high probability under the truncated Ising model, a sample $\bm{\sigma}$ has many neighboring configurations at Hamming distance 1, we construct an argument based on the Lovász Local Lemma (LLL), to guarantee the existence of a large number of satisfying assignments to $\Phi$ that differ from $\bm{\sigma}$ in exactly one bit. Using this powerful tool, however, requires control of the probabilities of partial spin assignments, which, given the tendency of the Ising model to contract into arbitrarily small portions of the hypercube and exhibit long-range correlations, can prove challenging. Counteracting this, our argument conditions on nodes outside of a specially crafted independent set $I$ of the graph $G$, which preserves the marginal distribution of any given spin $\mathbb{Pr}_{\beta^{*}}[\sigma_{i}|(\sigma_{1},...,\sigma_{i-1},\sigma_{i+1},...,\sigma_{n})]$ despite limited to a small fraction of the support of the measure, and collapses the Ising model into a product measure.
@@
We additionally note that the recent results DDDK (20) rely on sophisticated concentration inequalities derived from the fast mixing nature of Glauber dynamics on the Boolean hypercube and their relation to the Gibbs measure. In our model, these powerful tools are not applicable due to the fragmented nature of the truncation set, making the Glauber dynamics non-Ergodic; the inequalities only imply concentration within a connected component of $S$, which may be too small to be informative.
@@ 1.3 Related Work
The literature of parameter estimation in Markov Random Fields, and over hard-constrained/truncated measures, is vast. In light of this, we mention a brief collection of works relevant to our setting, and defer additional background and discussion to the appendix. Single sample estimation initiated by (Bes, 75; Cha, 07) has yielded a rich bounty of results ranging from the setting of the Ising model (Cha, 07; BM, 18; GM, 20; DDDK, 20), peer dependent logistic regression (DDP, 19; MNH+, 22; DDP, 20), higher order Ising models (MSB, 22), and robust inference over discrete distributions (DKSS, 21). BDNP (21) demonstrated the feasibility of single-sample learning in the context of the hard-core model, a size-weighted distribution over all independent sets in a graph $G$; following up on this, (GKK, 24; GGZ, 25) studied parameter inference in a product distribution truncated by the satisfying assignments of a $k$-SAT formula. The hard-constrained models studied in this work are a subset of the literature analyzing efficient parameter estimation and learning in truncated DGTZ (19, 18); FKT (22); DNS (23); NP (20) and censored distributions LMN (24); Ple (21); FKKT (21).
@@ 3 Learning Truncated Ising Models
In this section, we prove our main result, i.e., we provide a sufficient condition on the ”complexity” of the $k$-SAT formula $\Phi$, (and by extension the truncation set $S$) in terms of the size of the clauses $k$, the degree of the formula $d$ and the maximum degree of the graph $\Delta$ for efficient estimation of the inverse temperature parameter $\beta$. In advance of proving our result, we lay out some mild assumptions on the interaction matrix $A$ and Ising model $\mu_{G,\beta,S}$ which have been employed in past works GKK (24); GGZ (25); DDDK (20); DDP (19); Cha (07); BM (18).
@@
Let $\bm{\sigma}$ be a single sample from a truncated, $n$-dimensional Ising model satisfying Assumption 1. For all $k\geq\frac{4\Delta^{3}(1+\log(d^{2}k+1))}{\log(1+\exp(-2B))}$, if $n$ is sufficiently large, there exists an $\mathcal{O}(\Delta^{3}n\log(n))$-algorithm which takes as input $\bm{\sigma}$ and outputs an estimator $\hat{\beta}$ such that
@@
We compute $\hat{\beta}$ in time $\mathcal{O}(\Delta^{3}n\log n)$ by running projected gradient descent (PGD) on the normalized log-pseudolikelihood objective $n^{-1}\phi(\beta;\bm{\sigma})$. Standard results from convex optimization (e.g., BV (04)) imply that PGD converges to an $\epsilon$-optimal solution in $\mathcal{O}(\kappa\log(1/\epsilon))$ iterations, where $\kappa$ is the condition number of the objective, that is the ratio of the smoothness (i.e., the Lipschitz constant of the gradient) to the strong convexity parameter. In the sequel (Section 3.3), we demonstrate that the pseudo-likelihood objective is strongly convex with parameter $\Omega(n/\Delta^{3})$, implying the normalized objective is $\Omega(1/\Delta^{3})$-strongly convex. This, combined with an upper bound of $1$ on the norm of the normalized gradient of the log-pseudolikelihood, yields the condition number of the objective is $\kappa=\mathcal{O}(\Delta^{3})$. Due to the statistical limitations of the pseudolikelihood estimator, whose distance from the true parameter $\beta$ can be as large as $\mathcal{O}(\Delta^{3}/\sqrt{n})$, we set $\epsilon=1/\sqrt{n}$. Obtaining an accuracy of $\epsilon$, requires $\mathcal{O}(\Delta^{3}\log n)$ iterations, each requiring $\mathcal{O}(n)$ time, resulting in an overall runtime of $\mathcal{O}(\Delta^{3}n\log n)$.

### v2

# Learning the Inverse Temperature of Ising Models under Hard Constraints using One Sample

@@ 1 Introduction
Markov random fields (MRFs) are a common framework for analyzing high-dimensional distributions with complex conditional structures. A well-studied example of an MRF and the primary topic of inquiry in this paper is the Ising model ising1925beitrag, a probability measure $\mu_{G,\beta}$, over all assignments $\bm{\sigma}$ in the binary hypercube $\{-1,1\}^{n}$. The model is parameterized by a graph $G=(V,E)$ and inverse temperature $\beta$, taking the form $\mu_{G,\beta}\propto\exp\left(\beta\bm{\sigma}^{\top}A\bm{\sigma}\right)$, where $A$ being the adjacency matrix of $G$. The simplicity of the Ising model has led to widespread adoption in fields as disparate as statistical physics, finance, the social sciences, and computer vision, among others (see chatterjee2007estimation; qin2011polynomial; Harris (2013); Wang et al. (2015) and the references therein for a brief collection of examples). These applications have, in turn, motivated a substantial body of research on efficient sampling Bresler (2015); lubetzky2013cutoff; sly2012computational, rigorous testing Daskalakis et al. (2019a), and principled inference of the inverse temperature parameter and interaction matrix Dagan et al. (2020, 2021) under the framework of the Ising model.
Beginning from the work of chatterjee2007estimation, there has been substantial interest in the task of estimating inverse temperature parameter $\beta$, given only one sample $\bm{\sigma}\sim\mu_{G,\beta}$ and the graph $G$, using the maximum pseudolikelihood estimator of Besag (1975). This setup subsumes the setting of multi-sample estimation as $\ell$ samples of an Ising model over $n$ nodes is equivalently a single sample of an $n\ell$ Ising model over $\ell$ disconnected components of a graph. This line of work is driven by the inherent technical constraints of network data, wherein it is often impractical to obtain independent observations of the same network responses daskalakis2019regression; Dagan et al. (2020). Interestingly, despite both static and dynamic phase transitions in model behavior as $\beta$ varies which can render sampling computationally intractable (NP-hard) GALANIS_2016, it remains possible to construct consistent estimators of $\beta$ and the interaction matrix $A$, provided that $\beta=\mathcal{O}(1)$ Dagan et al. (2020, 2021); daskalakis2019regression; Mukherjee et al. (2022a).
@@
In many real-world applications, however, we face not only soft constraints, which influence the model behavior by introducing correlations or dependencies, while the full support of the measure remains intact, but also hard constraints: Certain configurations are outright forbidden, and entire regions of the configuration space are excluded from the support of the distribution. Such hard-constrained models (also called truncated) arise naturally in applications involving high-dimensional, interconnected systems with strict feasibility requirements. One particularly notable example of which arises in the context of spatial transcriptomics—a biological framework for characterizing gene and protein expression in cells within organic tissue, relative to their spatial organization. The local relationships between cells are often represented as nodes in a graph with edges linking cells together that are close in physical space (Eng et al., 2019). Associated with each node is a set of measurements of expression to capture how the relationships between cells impact the phenotype of a given node. The complex relationship between genes often forces certain configurations of expressions to be infeasible. In fact, the phenomenon of lateral inhibition can cause a cell expressing gene one to prevent its neighbors from expressing it as well, instead causing them to express gene two, as seen in the Notch-Delta pathway Ghosh and Tomlin (2001). The hard constraints discussed in the above setting are not unique to spatial transcriptomics and are also commonly found in the context of channel assignments in communication networks (Zafer and Modiano, 2006), carrier-sense multiple access networks (Durvy and Thiran, 2006; Durvy et al., 2009), and multicasting networks (Karvo et al., 2002; Luen et al., 2006) among others.
In this work, we study the problem of parameter estimation in $n$-dimensional Ising models that are hard-constrained to the satisfying assignments of a bounded-degree $k$-SAT formula $\Phi$ expressed in CNF (Conjunctive Normal Form), using one sample. This means that we have access to a sample from an Ising model, conditioned that it only takes values in a subset $S\subset\{\pm 1\}^{n}$ that is represented through the satisfying assignments of a $k$-SAT formula, adopting the framework from galanis2024learning. The choice of the $k$-SAT framework is motivated by the fundamental observation that any truncation set $S\subseteq\{\pm 1\}^{n}$ can be exactly realized as the solution space of such a formula, provided the number of clauses is sufficiently large. This representation is particularly advantageous for statistical inference, as the structural parameters, in particular the clause size $k$ and the variable degree, serve as natural controls for the complexity of the induced correlations in the analysis of the Maximum Pseudolikelihood Estimator (MPLE).
@@
More generally, learning in truncated MRFs using one or multiple samples has been studied in the context of discrete product distributions truncated by the set of satisfying assignments of $k$-SAT formulas galanis2024learning; Galanis et al. (2025), more generally by truncated sets with combinatorial structure Fotakis et al. (2022), the hard-core model, and integer valued spins constrained over proper $H$-colorings blanca2018structure; bhattacharya2021parameterestimationundirectedgraphical. Our key deviation from the aforementioned works stems from the fact that the Ising model is not a product distribution, and common tools used to control the concentration of measure on the hypercube do not apply. This, moreover, induces two sources of interdependence, namely from the model itself and from the structure of the truncation set. With this background in mind, we seek to address the following challenge.
@@ 1.1 Our Results
Our main contribution is an affirmative answer to the previous challenge, by providing a sufficient condition on the $k$-SAT formula that induces the truncation set, in terms of the maximum degree $\Delta$ of $G$. We begin by formally defining the class of truncated Ising measures that is the primary inquiry of this work. Given a graph $G$, with associated adjacency matrix $A$, and inverse temperature $\beta$ we define the pmf of a truncated Ising model for any $\bm{\sigma}\in\{\pm 1\}^{n}$ to be
@@
where $\mathbf{1}\{\bm{\sigma}\in S\}$ captures the indicator function determining if $\bm{\sigma}\in S$ and $Z_{\beta,S}$ is a renormalization term denoted the partition function. In our case, $S$ is expressed as the set of satisfying assignments of a bounded degree $k$-SAT formula $\Phi_{n,k,d}=\Phi$ in constrained normal form. Formally, let $\Phi$ be a formula in conjunctive normal form over variables $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ and clause set $\mathcal{C}$, subject to the structural constraints that each clause contains exactly $k$ literals (variables or their negations) and each variable occurs in at most $d$ clauses. We identify each configuration $\bm{\sigma}\in\{-1,1\}^{n}$ with a truth assignment via the mapping where variable $v_{i}$ is true if $\sigma_{i}=1$ and false if $\sigma_{i}=-1$. A configuration $\bm{\sigma}$ is then said to satisfy $\Phi$ if and only if every clause $C\in\mathcal{C}$ contains at least one literal that evaluates to true under this assignment. Recall, any subset of the hypercube can be represented as the set of satisfying assignments of a degree-$d$ $k$-SAT formula, provided that $d$ is sufficiently large.
@@
Remark (Consistency).
Notice when $\Delta$ is $\mathcal{O}(1)$, our estimate achieves $\mathcal{O}(1/\sqrt{n})$-consistency, matching the minimax rate for parameter estimation. Likewise, the restriction of $\Delta$ to be on the order of $o(n^{1/6})$ ensures the error of our estimator $\hat{\beta}$ is asymptotically diminishing and in turn consistent.
Remark (Lower Bounds).
We further note that Galanis et al. galanis2024learning recently established a lower bound for learning discrete Boolean distributions, demonstrating that estimation is information-theoretically impossible when $k\leq\log(d)-\log(k)+\Theta(1)$. This result relies on constructing a $k$-SAT instance with a single satisfying assignment. Importantly, this lower bound extends to our setting, implying that our results are optimal up to a constant factor when $G$ is a graph of bounded degree. While subsequent work Galanis et al. (2025) has provided a tighter bound, those techniques do not immediately translate to our context; the quadratic Hamiltonian induces a rugged energy landscape distinct from boolean product setting, necessitating different analytical tools.
@@ 1.2 Technical Overview
Given a single-parameter exponential family like we focus on, a natural approach to estimating the parameter is to find the maximum likelihood estimate. However, the computational intractability of the partition function $Z_{\beta,S}$ for Ising models (see GALANIS_2016 and the references therein) renders this approach infeasible. In light of these challenges, we utilize the maximum pseudolikelihood estimator introduced by Besag (1975) and provided below.
@@
We note that the second equality holds due to the monotonicity of the $\log(\cdot)$ function. Towards demonstrating the consistency of the maximum (log)-pseudolikelihood estimate $\hat{\beta}$, we follow the first and second derivative paradigm outlined by Chatterjee chatterjee2007estimation; daskalakis2019regression; galanis2024learning; Galanis et al. (2025), which involves showing,
@@
The first condition ensures the derivative of the log-pseudolikelihood objective with respect to the true model parameters $\beta$ divided by $n$ is close to 0, which is the value of the gradient of $\phi$ computed at the estimator, which in turn implies $\beta$ is an approximate stationary point of the objective. Moreover, by demonstrating that the second derivative of the objective $\nabla^{2}_{\beta}\phi(\beta;A,\bm{\sigma})$ is $\Omega(n/\Delta^{3})$-strongly convex with probability 1o(1)1-o(1) over a draw of the truncated Ising model, it implies that approximate stationary points of the objective are close in Euclidean distance to the optimum. These two facts combine to show the proximity of the optimum of the log-pseudolikelihood objective to $\beta$.
@@
Showing both of these conditions hold simultaneously is made complex due to the highly non-uniform measure induced by conditional dependencies of both the interaction matrix $A$ and the truncation set $S$. To demonstrate the first condition, we craft upper bounds on the variance of the first derivative of $\phi$, using the technique of exchangeable pairs pioneered by chatterjee2007estimation, which, when combined with Chebyshev’s inequality, implies an upper bound in probability. The primary challenge of this work lies in establishing the second condition. Previous works which used the deterministic structure of the interaction matrix to guarantee the concavity of the objective. By contrast, in our setting, the second derivative is governed by the local geometry of the truncation set $S$ in the vicinity of the sample $\bm{\sigma}$. The Hessian $\nabla_{\beta}^{2}$ is computed by summing a function of the local fields $m_{i}(\bm{\sigma})=\langle A_{i},\bm{\sigma}\rangle$, where the support of this sum is limited to the valid neighbors of $\bm{\sigma}$ in $S$ with respect to the Hamming distance  [^1] The Hamming distance metric counts the number of differing indices between two vectors. Thus, a neighbor at Hamming distance one is obtained by flipping exactly one index of $\bm{\sigma}$.
@@
To show that, with high probability under the truncated Ising model, a sample $\bm{\sigma}$ has many neighboring configurations at Hamming distance 1, we construct an argument based on the Lovász Local Lemma (LLL), to guarantee the existence of a large number of satisfying assignments to $\Phi$ that differ from $\bm{\sigma}$ in exactly one bit. Using this powerful tool, however, requires control of the probabilities of partial spin assignments, which, given the tendency of the Ising model to contract into arbitrarily small portions of the hypercube and exhibit long-range correlations, can prove challenging. Counteracting this, our argument conditions on nodes outside of a specially crafted independent set $I$ of the graph $G$, which preserves the marginal distribution of any given spin $\mathbb{Pr}_{\beta^{*}}[\sigma_{i}|(\sigma_{1},...,\sigma_{i-1},\sigma_{i+1},...,\sigma_{n})]$ despite limited to a small fraction of the support of the measure, and collapses the Ising model into a product measure. To conclude the strong convexity of $\phi(\cdot;A,\bm{\sigma})$ it remains to show a lower bound on $m_{i}(\bm{\sigma})$ in probability via a coupling argument which exploits the underlying edge structure of the connectivity graph $G$.
@@
We additionally note that the recent results on estimating Ising models using the pseudolikelihood approach Dagan et al. (2020) rely on sophisticated concentration inequalities derived from the fast mixing nature of Glauber dynamics on the Boolean hypercube and their relation to the Gibbs measure in order to conclude the strong convexity of the PL objective and bound the gradient with high probability. In our model, these powerful tools are not applicable due to the fragmented nature of the truncation set, rendering the domain of our measure into disconnected islands and make the Glauber dynamics non-Ergodic; the inequalities only imply concentration within a connected component of $S$, which may be too small to be informative. We lastly emphasize that our polynomial time algorithm does not scale with the size of the truncation set, enabling inference despite $S$ being small with regards to the entire hypercube.
@@ 1.3 Related Work
The literature of parameter estimation in Markov Random Fields, and over hard-constrained and truncated measures, is vast. In light of this, we mention a brief collection of works relevant to our setting, and defer additional background and discussion to the appendix. Single sample estimation initiated by (Besag, 1975; chatterjee2007estimation) has yielded a rich bounty of results ranging from the setting of the Ising model (chatterjee2007estimation; Bhattacharya and Mukherjee, 2018; Ghosal and Mukherjee, 2020; Dagan et al., 2020), peer dependent logistic regression (daskalakis2019regression; Mukherjee et al., 2022a; Daskalakis et al., 2020), higher order Ising models (Mukherjee et al., 2022b), and robust inference over discrete distributions (Diakonikolas et al., 2021). bhattacharyya2021efficient demonstrated the feasibility of single-sample learning in the context of the hard-core model, a size-weighted distribution over all independent sets in a graph $G$; following up on this, (galanis2024learning; Galanis et al., 2025) studied parameter inference in a product distribution truncated by the satisfying assignments of a $k$-SAT formula. The hard-constrained models studied in this work are a subset of the literature analyzing efficient parameter estimation and learning in truncated Daskalakis et al. (2019b, 2018); Fotakis et al. (2022); De et al. (2023); Nagarajan and Panageas (2020) and censored distributions Lugosi et al. (2024); Plevrakis (2021); Fotakis et al. (2021).
@@ 3 Learning Truncated Ising Models
In this section, we prove our main result, i.e., we provide a sufficient condition on the ”complexity” of the $k$-SAT formula $\Phi$, (and by extension the truncation set $S$) in terms of the size of the clauses $k$, the degree of the formula $d$ and the maximum degree of the graph $\Delta$ for efficient estimation of the inverse temperature parameter $\beta$. In advance of proving our result, we lay out some mild assumptions on the interaction matrix $A$ and Ising model $\mu_{G,\beta,S}$ which have been employed in past works galanis2024learning; Galanis et al. (2025); Dagan et al. (2020); daskalakis2019regression; chatterjee2007estimation; Bhattacharya and Mukherjee (2018).
@@
Remark (Assumptions).
The assumption of graph connectivity ensures that the interaction matrix contains sufficient signal energy for consistent parameter estimation. This requirement plays a role analogous to the lower bounds on the Frobenius norm of the interaction matrix posited in chatterjee2007estimation; daskalakis2019regression; Bhattacharya and Mukherjee (2018). Absent such a structural guarantee, the log-partition function $Z_{\beta,S}$ may fail to diverge asymptotically with $n$, which prevents the pseudolikelihood objective from achieving consistency. This pathology is exemplified by the Curie-Weiss model with couplings scaling as $1/n$; there, the Frobenius norm of the interaction matrix is $\mathcal{O}(1)$. In contrast, our connectivity and degree assumptions imply that the squared Frobenius norm of $A$ scales as $\Omega(n/\Delta^{2})$, which diverges to infinity, thereby ensuring the problem is well-posed.
While general specifications of the Ising model allow for arbitrary coupling strengths $A_{ij}\in\mathbb{R}$, we restrict the interaction magnitudes to a uniform value $|A_{ij}|=1/\Delta$ towards isolating the structural component of peer influence. This constraint posits that the social pressure exerted by any single neighbor is functionally equivalent, thereby modeling a “peer effect” Bertrand et al. (2000); Sacerdote (2001); Duflo and Saez (2003) where individual decisions are driven by the collective consensus of the local group rather than the idiosyncratic intensity of specific dyadic relationships. By imposing this homogeneity, we ensure that the model captures the emergent coordination arising from network topology and group alignment, rather than being dominated by outliers with arbitrarily strong pairwise connections.
@@
Let $\bm{\sigma}$ be a single sample from a truncated, $n$-dimensional Ising model satisfying Assumption 1. For all $k\geq\frac{4\Delta^{3}(1+\log(d^{2}k+1))}{\log(1+\exp(-2B))}$, if $n$ is sufficiently large, there exists an $\mathcal{O}(\Delta^{3}n\log n)$-time algorithm which takes as input $\bm{\sigma}$ and outputs an estimator $\hat{\beta}$ such that
@@
We compute $\hat{\beta}$ in time $\mathcal{O}(\Delta^{3}n\log n)$ by running projected gradient descent (PGD) on the normalized log-pseudolikelihood objective $n^{-1}\phi(\beta;\bm{\sigma})$. Standard results from convex optimization (e.g., BV04) imply that PGD converges to an $\epsilon$-optimal solution in $\mathcal{O}(\kappa\log(1/\epsilon))$ iterations, where $\kappa$ is the condition number of the objective, that is the ratio of the smoothness (i.e., the Lipschitz constant of the gradient) to the strong convexity parameter. In the sequel (Section 3.3), we demonstrate that the pseudo-likelihood objective is strongly convex with parameter $\Omega(n/\Delta^{3})$, implying the normalized objective is $\Omega(1/\Delta^{3})$-strongly convex. This, combined with an upper bound of $1$ on the norm of the normalized gradient of the log-pseudolikelihood, yields the condition number of the objective is $\kappa=\mathcal{O}(\Delta^{3})$. Due to the statistical limitations of the pseudolikelihood estimator, whose distance from the true parameter $\beta$ can be as large as $\mathcal{O}(\Delta^{3}/\sqrt{n})$, we set $\epsilon=1/\sqrt{n}$. Obtaining an accuracy of $\epsilon$, requires $\mathcal{O}(\Delta^{3}\log n)$ iterations, each requiring $\mathcal{O}(n)$ time, resulting in an overall runtime of $\mathcal{O}(\Delta^{3}n\log n)$.